# Two-Player Zero-Sum Games with Bandit Feedback

### v2

@@ 6 ETC-TPZSG-AE-NUE Algorithm
In this section, we extend the adaptive elimination algorithm by introducing a variant, ETC-TPZSG-AE-NUE. It uses a non-uniform exploration method The non-uniform exploration in these algorithms is based on players’ strategies, which indirectly reflect the probability of being NE. The key difference from the previous algorithm is that Algorithm 3 explores action pairs in proportion to the strategies of the players derived from the estimated payoff matrix. Thus, it is employed to reduce the number of plays for suboptimal action pairs more effectively.
@@
The estimated strategies of the players serve as a basis to determine the exploration time in our proposed algorithms. Furthermore, the probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}_{i,t}\hat{q}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$. Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
We incorporate the strategies of the players $\hat{p}_{t}$ and $\hat{q}_{t}$ together with $k_{t}$. Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
$\begin{aligned} &\\ &\hat{P}_{t}=\hat{p}_{t}\hat{q}_{t}\\ &\hat{p}_{t}=\max_{p}\min_{q}p^{T}\hat{A}q\\ &\hat{q}_{t}=\min_{q}\max_{p}p^{T}\hat{A}q\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (24)
@@
1:Input:
@@
3:$m$: number of actions for row player
4:$l$: number of actions for column player
5:$T$: time horizon $\triangleright$ $1\leq ml\ll T$
[…]
7:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$, $\hat{A}=[0]^{m\times l}$ and $P_{ij,0}=\frac{1}{ml},\forall i,j$
8:In round $t=0,1,2,...,\left\lfloor\frac{1}{4}\log_{2}\frac{T}{e}\right\rfloor$;
9:    Action Pair Selection:
10:     If $\mathbf{|S_{t}|>1}:$ $\triangleright$ Exploration Phase
11:      Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (25)
12:      Update $\hat{A}$ using (4).
13:     Else: $\triangleright$ Committing Phase
14:      Play with a unique action pair in $S_{t}$ until step $T$
15:    Action Pair Elimination:
16:     Remove the action pairs $(i,j)$ NOT satisfying the following for $\varepsilon_{t}=\sqrt{\frac{4\sigma^{2}}{k_{t}}\ln{(\frac{\hat{\Delta}_{t}^{2}T}{16\sigma^{2}})}}$:
@@
17:    Update:
18:     $\hat{P}_{ij,t}$ by solving the linear system in (24).
19:     $S_{t+1}$ by "Action Pair Elimination" step
20:     $\hat{\Delta}_{t+1}=\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
Algorithm 3 ETC-TPZSG-AE-NUE
@@
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}+\Delta_{ij}^{*}T\Big)$ (26)
@@ 7 Experiments
In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], the only existing method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs. We use a $2\times 2$ game matrix by Gaussian payoffs with variance $\sigma^{2}$ and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list. We note that a smaller $\sigma$ leads to a smaller $\Delta$.
@@ Acknowledgments.
Appendix A Subgaussian Properties
This section collects some supplementary results from [20] that are used repeatedly throughout our proofs. These results concern standard properties of subgaussian random variables such as tail inequalities. They are included here for completeness and easy reference as they play a critical role in our analyses. The statements are not original, we include only the parts that are directly useful for our purposes.
Theorem 5 ([20, Theorem 5.3]).
@@ Appendix B Proof of Theorem 1
To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the committing phase. During exploration, each action pair is played a fixed number of times $k$ to estimate their mean rewards, potentially incurring regret if suboptimal action pairs are played. Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one. Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it as a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
@@
$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (29)
$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}))\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))}$ (30)
$=\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)}$ (31)
$\leq\sqrt{\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (32)
@@
where $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
To get (30), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, that is, for any collection of events $e_{1},e_{2},\ldots,e_{n}$, we have $\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i})$ for all $i\in{1,\ldots,n}$. The inequality (32) holds by using the followings:
$\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)=\mathbb{P}(\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})\leq A(i,j^{\prime})-A(i^{\prime},j^{\prime}))$ (38)
$\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)=\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)\leq A(i^{\prime},j^{\prime})-A(i^{\prime},j))$ (39)
Then, we apply $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}\geq A(i,j^{\prime})-A(i^{\prime},j^{\prime}),\Delta_{i^{\prime}j^{\prime}}^{\text{min}}\geq A(i^{\prime},j^{\prime})-A(i^{\prime},j)$, which are true from their definitions. For step (33), we apply Theorem 5 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable. At this point, we need to show that the differences $\hat{A}(i,j^{\prime})+A(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})$ and $\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)-A(i^{\prime},j)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian but it is enough to check that for any action pairs $(i,j)$ and $(x,y)$ in $S_{A}$, $\hat{A}(i,j)-\hat{A}(x,y)$ are $\sqrt{2\sigma^{2}/k}$-subgaussian. To establish this, we utilize Lemma 1, which provides key properties of subgaussian random variables. Specifically, both $\hat{A}(i,j)$ and $\hat{A}(x,y)$ are $\sqrt{\sigma^{2}/k}$-subgaussian as the average payoffs are computed by the equation (4) and during the exploration phase, each action pair is played $k$ times. By the lemma, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$, which follows that $\hat{A}(i,j)-\hat{A}(x,y)$ is $\sqrt{2\sigma^{2}/k}$-subgaussian for any action pairs $(i,j)$ and $(x,y)$. To get the last inequality (35), we observe that
$\frac{\Delta_{i^{\prime}j^{\prime}}^{2}}{2}=\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}\leq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$
where the inequality follows from the fact that $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}}-\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq 0$.
As we mention before, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE. It involves an inequality based on the minimum of individual probabilities. Similarly, let follow this:
@@
$\leq\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))$ (41)
$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})),\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j))\}$ (42)
$=\min\{\mathbb{P}(\hat{A}(i,j^{\prime})-\hat{A}(i^{\prime},j^{\prime})\leq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i^{\prime},j)\leq 0)\}$ (43)
@@
where to get the inequality (46), we use the following:
$\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}\geq\frac{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}+(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}}{2}$ (49)
If we assume that their maximum is $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}$, then it will imply $(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$. Similarly, if $\max\{(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2},(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\}=(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}$, it will give us $(\Delta_{i^{\prime}j^{\prime}}^{\text{min}})^{2}\geq(\Delta_{i^{\prime}j^{\prime}}^{\text{max}})^{2}$. They are true by assumptions, then, (49) holds.
Thus, both approaches provide the same upper bound for the probability of misidentified NE. In the regret analysis, we will focus on the loss due to playing with suboptimal (or not true NE) action pairs. Thus, after $T$ times playing, $1\leq Nk\leq T$ where $N=ml$ as the number of action pairs, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$
where $\mathbb{E}[n_{ij,T}]$ is the expected number of playing the action pair $(i,j)$. We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration. After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of true one $(i^{*},j^{*})$ because we expect the players to select NE action pair as optimal after the exploration phase. Since we already have the probability of playing a misidentified NE from (37) or (48), we can write
$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$

### v3

@@ 6 ETC-TPZSG-AE-NUE Algorithm
In this section, we extend the adaptive elimination algorithm by introducing a variant, ETC-TPZSG-AE-NUE. It uses a non-uniform exploration method, which is based on players’ strategies, which indirectly reflect the probability of being NE. The key difference from the previous algorithm is that Algorithm 3 explores action pairs in proportion to the strategies of the players derived from the estimated payoff matrix. Thus, it is employed to reduce the number of plays for suboptimal action pairs more effectively.
@@
The estimated strategies of the players serve as a basis to determine the exploration time in Algorithm 3. The probability of selecting an action pair $(i,j)$ is estimated by $\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}$, which satisfies $\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=1$, where $(\hat{p}_{t}^{*},\hat{q}_{t}^{*})$ is a mixed NE in round $t$. In other words, once optimal strategies are determined, we calculate the joint probability $\hat{P}_{ij,t}$, which represents the probability that the row player chooses action $i$ and the column player chooses action $j$ in round $t$. Consequently, in Algorithm 3, the exploration time for each action pair $(i,j)$ is proportional to $\hat{P}_{ij,t}$, allowing action pairs with higher estimated probabilities to be played more frequently.
We incorporate the strategies of the players $\hat{p}^{*}_{t}$ and $\hat{q}^{*}_{t}$ together with $k_{t}$. Specifically, in each round $t$, this involves estimating the optimal strategies for each player by solving a linear system using the estimated payoff matrix $\hat{A}$ as follows:
$\begin{aligned} &\\ &\hat{P}_{ij,t}=\hat{p}^{*}_{i,t}\hat{q}^{*}_{j,t}\\ &\hat{p}^{*}_{t}=\operatorname*{argmax}_{\hat{p}}\min_{\hat{q}}\hat{p}^{T}\hat{A}\hat{q}\\ &\hat{q}^{*}_{t}=\operatorname*{argmin}_{\hat{q}}\max_{\hat{p}}\hat{p}^{T}\hat{A}\hat{q}\\ \end{aligned}\qquad\begin{aligned} &\text{Subject to:}\\ &\sum_{i\in S_{x}}\hat{p}_{i,t}=1,\quad\hat{p}_{i,t}\geq 0,\ \forall i\in S_{x}\\ &\sum_{j\in S_{y}}\hat{q}_{j,t}=1,\quad\hat{q}_{j,t}\geq 0,\ \forall j\in S_{y}\\ &\sum_{(i,j)\in S_{A}}\hat{P}_{ij,t}=\sum_{i,j}\hat{p}_{i,t}\hat{q}_{j,t}=1\end{aligned}$ (25)
@@
Algorithm 3 ETC-TPZSG-AE-NUE
1:Input: $S_{A},m,l,\sigma,T(1\leq ml\ll T)$,
2:Initialize: $\hat{\Delta}_{0}=4\sigma$, $S_{0}=S_{A}$ and $\hat{A}=[0]^{m\times l}$
3:for $t=0,1,2,...,\left\lfloor\frac{1}{2}\log_{2}\frac{T}{e}\right\rfloor$ do
4:  if $\mathbf{|S_{t}|>1}:$ then $\triangleright$ Exploration Phase
5:   Explore each $(i,j)$ in $S_{t}$ $k_{ij,t}$ times, defined in (26)
6:   Update $\hat{A}$ using (4)
@@
8:   Play with a unique action pair in $S_{t}$ until step $T$
9:  end if
10:  Remove the action pairs $(i,j)$ from $S_{t}$ NOT satisfying the following for $\varepsilon_{t}$ in (22): $\triangleright$ Elimination Phase
@@
11:  Solve the linear system in (25) to get $\hat{P}_{ij,t+1}$
12:  $S_{t+1}\leftarrow S_{t}$
13:  $\hat{\Delta}_{t+1}\leftarrow\frac{\hat{\Delta}_{t}}{2}$ where $\hat{\Delta}_{t}=2^{-t+2}\sigma$
14:end for
@@
$R_{T}^{*}\leq\sum_{(i,j)\in S_{A_{1}}}\Delta_{ij}^{*}\Big(1+\frac{768\sigma^{2}}{\Delta_{ij}^{2}}+\frac{512\sigma^{2}}{\Delta_{ij}^{2}}\ln{\Big(\frac{\Delta_{ij}^{2}T}{256\sigma^{2}}\Big)}\Big)+\sum_{(i,j)\in S_{A_{2}}}\Big(\Delta_{ij}^{*}\frac{512\sigma^{2}}{\lambda^{2}}\Big)+\max_{(i,j)\in S_{A_{2}}}\Delta_{ij}^{*}T$ (27)
@@ 7 Experiments
In Figures 3 and 3, we compare the proposed algorithms and Tsallis-INF [18], very recent method with instance-dependent bounds for TPZSGs, based on cumulative regret, computed as the absolute difference between the game value and the payoff from the action played to avoid negative values, averaged over $10^{2}$ runs. We use a $2\times 2$ game matrix by Gaussian payoffs with variance $\sigma^{2}$ and $k$ exploration rounds per action pair for ETC-TPZSG is randomly selected from a predefined list. We note that a smaller $\sigma$ leads to a smaller $\Delta$.
@@ Acknowledgments.
Appendix A Useful Properties
In this section, we present several properties that are used throughout the proofs. Moreover, this section collects some supplementary results from [20]. These results concern the standard properties of subgaussian random variables. They are included here for completeness and easy reference, as they play a critical role in our regret analyses.
Lemma 1.
For any $a,b>0$, we can write the following:
$\max\{a^{2},b^{2}\}\geq\frac{a^{2}+b^{2}}{2}$ (28)
Proof.
If we assume that their maximum is $a^{2}$, then we have $a^{2}\geq a^{2}+b^{2}/2$, which implies $a^{2}\geq b^{2}$. Similarly, if $\max\{a^{2},b^{2}\}=b^{2}$, it gives us $b^{2}\geq a^{2}$. They are true by assumptions, and thus we conclude the proof. ∎
Lemma 2.
For any $a,b>0$, the following holds:
$a^{2}+b^{2}\geq\frac{(a+b)^{2}}{2}$ (29)
Proof.
The proof follows from the fact that $(a-b)^{2}\geq 0$. ∎
Lemma 3.
For events $e_{1},e_{2},\cdots,e_{n}$, we have the following for any $i\in\{1,2,\cdots,n\}$:
$\mathbb{P}(e_{1}\cap e_{2}\cap\cdots\cap e_{n})\leq\mathbb{P}(e_{i}).$ (30)
Proof.
For any $i\in\{1,2,\cdots,n\}$, we observe that $e_{1}\cap e_{2}\cap\cdots\cap e_{n}\subseteq e_{i}$. Thus, the proof comes from the fact that an intersection of events cannot be more likely than any of the events involved. ∎
Lemma 4 ([20, Theorem 5.3]).
@@ Appendix B Proof of Theorem 1
To analyze the Nash regret of the Algorithm 1, we begin by decomposing the total expected regret into two phases: the exploration phase and the commit phase. During exploration, each action pair is played a fixed number of times $k$ to estimate their mean payoffs, potentially incurring regret if suboptimal action pairs are played. Once the algorithm commits to the empirically best action pair, which is the pure NE, the regret accumulates only if this action pair is not the true optimal one. Our goal is to bound the expected regret by quantifying the probability of committing to a suboptimal action pair, which we can call it a misidentified NE, based on the estimates of their payoffs obtained during the exploration phase.
@@
$\leq\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))$ (33)
$\leq\sqrt{\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime}))\ \mathbb{P}(\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))}$ (34)
$=\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})\geq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})\geq 0)}$ (35)
$\leq\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-(\hat{A}(i_{1},j^{\prime})-A(i_{1},j^{\prime}))\geq\Delta_{i^{\prime}j^{\prime}}^{\text{max}})}$
$\quad\times\sqrt{\mathbb{P}(\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-(\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime}))\geq\Delta_{i^{\prime}j^{\prime}}^{\text{min}})}$ (36)
@@
where we assume that $A(i_{1},j^{\prime})=\max_{i\in S_{x}}A(i,j^{\prime})$ and $A(i^{\prime},j_{1})=\min_{i\in S_{x}}A(i^{\prime},j)$. On the other hand, we note $\Delta_{i^{\prime}j^{\prime}}^{\text{max}}=\max_{i}A(i,j^{\prime})-A(i^{\prime},j^{\prime})$, $\Delta_{i^{\prime}j^{\prime}}^{\text{min}}=A(i^{\prime},j^{\prime})-\min_{j}A(i^{\prime},j)$ and $\Delta_{i^{\prime}j^{\prime}}=\Delta_{i^{\prime}j^{\prime}}^{\text{max}}+\Delta_{i^{\prime}j^{\prime}}^{\text{min}}$.
To obtain (33), we utilize the fact that the probability of the intersection of multiple events is at most that of any individual event, as stated in 3. For step (37), we apply Lemma 4 with $\sqrt{2\sigma^{2}/k}$-subgaussian random variable. At this point, we need to show that the differences $\hat{A}(i^{\prime},j^{\prime})-A(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})+A(i_{1},j^{\prime})$ and $\hat{A}(i^{\prime},j_{1})-A(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})+A(i^{\prime},j^{\prime})$ are $\sqrt{2\sigma^{2}/k}$-subgaussian. To establish this, we utilize Lemma 5, which provides key properties of subgaussian random variables. Specifically, for any $(i,j)\in S_{A}$, $\hat{A}(i,j)-A(i,j)$ are $\sqrt{\sigma^{2}/k}$-subgaussian since during the exploration phase, each action pair is played $k$ times. Then by Lemma 5, the difference of two independent subgaussian random variables with parameters $\sqrt{\sigma^{2}/k}$ is subgaussian with parameter $\sqrt{2\sigma^{2}/k}$. To obtain the last inequality (39), we apply Lemma 2.
As we mentioned earlier, there is an alternative approach to bound $\mathbb{P}(E_{1}\cap E_{2})$ to find a bound for the probability of a misidentified NE. It involves an inequality based on the minimum of individual probabilities. Similarly, let follow this:
@@
$\leq\mathbb{P}(\hat{A}(i_{1},j^{\prime})\leq\hat{A}(i^{\prime},j^{\prime})\wedge\hat{A}(i^{\prime},j^{\prime})\leq\hat{A}(i^{\prime},j_{1}))$ (43)
$\leq\min\{\mathbb{P}(\hat{A}(i^{\prime},j^{\prime})-\hat{A}(i_{1},j^{\prime})\geq 0)\ \mathbb{P}(\hat{A}(i^{\prime},j_{1})-\hat{A}(i^{\prime},j^{\prime})\geq 0)\}$ (44)
@@
where to get the inequality (47), we use Lemma 1.
Thus, both approaches provide the same upper bound for the probability of misidentified NE. In the regret analysis, we will focus on the loss due to playing with suboptimal (or not true NE) action pairs. Thus, after playing $T$ times, $1\leq Nk\leq T$ where $N=ml$ is the number of action pairs, we can write the Nash regret as
$R_{T}^{*}=\sum_{(i,j)\in S_{A}}\Delta_{ij}^{*}\mathbb{E}[n_{ij,T}]$ (50)
where $\mathbb{E}[n_{ij,T}]$ is the expected number of plays for the action pair $(i,j)$. We already know that each action pair $(i,j)\in S_{A}$ is played at least $k$ times for exploration. After the exploration step, we will consider playing the misidentified NE $(i^{\prime},j^{\prime})$ instead of the true one $(i^{*},j^{*})$ because we expect the players to select the NE action pair as optimal after the exploration phase. Since we already have the probability of playing a misidentified NE from (41) or (49), we can write
$\mathbb{E}[n_{ij,T}]\leq k+(T-Nk)\exp{\Big(-\frac{k\Delta_{ij}^{2}}{16\sigma^{2}}\Big)}.$ (51)